OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 4 7\8

rebeccaxhawaii (rebeccaxhawaii):

whats 4*8+? then add 7

rebeccaxhawaii (rebeccaxhawaii):

then put that number over 8

OpenStudy (anonymous):

oh i get it now thanks

rebeccaxhawaii (rebeccaxhawaii):

no problem c:

OpenStudy (skullpatrol):

$$\Huge 4 \dfrac{7}{8} = 4 + \dfrac{7}{8} = \dfrac{4}{1} + \dfrac{7}{8} =$$ $$\Huge = \dfrac{4•8}{1•8} + \dfrac{7}{8} = \dfrac{4•8 +7}{8}$$